OpenStudy (anonymous):

Find tan θ if sec θ = square root of thirty seven divided by six and sin θ < 0. negative square root of thirty seven negative one divided by six -6 6

OpenStudy (gorv):

\[\tan \theta=\sqrt{\sec^2 \theta +1}\]